### 19 Mar 2026 v2

# Designing a low-loss high reflectivity mirror for gravitational waves detectors by combining a dielectric metasurface and multilayer stacks

Abstract
Future generations of gravitational wave detectors require increased sensitivity, for which the availability of large mirrors with high reflectivity and low mechanical loss is essential. Current amorphous multilayer mirror designs present constraining limitations in terms of thermal noise. These mirrors require a large number of thin film layers to achieve near-perfect reflectivity. However, the thermal noise generated by this type of stack increases with the number of layers used. Reducing thermal noise is therefore very challenging and highlights the need for new technical solutions that can address this specific issue. Here, we provide insights into the expected performance of mirrors that combine a resonant metasurface with a multilayer stack. The suggested mirror design ensures the high reflectivity required for interferometric gravitational wave detectors, while using fewer layers of properly selected materials. It allows to reduce the total thickness of the material with the poorest thermal-noise performance, namely TiO2:Ta2O5, by a factor of more than 3, making it a promising option for potentially reducing thermal noise as well.
I Introduction
Since late 2015, the scientific community has demonstrated its ability to detect gravitational waves [1] using terrestrial observatories based on giant Michelson interferometers equipped with extreme-performance end mirrors [2]. These mirrors are made of high-quality, thick silica substrates coated with an all-dielectric multilayer stack that increases surface reflectivity to nearly one (typically 99.9995%). Today, however, the thermal noise of these amorphous dielectric stacks is the limiting factor of the interferometric detectors in the 50-300 Hz frequency range. The two dielectric materials used in all the most mirrors manufactured thus far are titania-doped tantala (TiO2:Ta2O5) as the high-index material and silica (SiO2) as the low-index material, both deposited using ion beam sputtering (IBS), because such thin-film materials offer the best trade-off between high refractive index contrast, low optical absorption, low thermal noise and scalability to large size [3]. To improve the thermal noise performance of these coatings, a wide range of research has been conducted, a popular option being optimizing the formula of the reflective stacks to avoid using the material with the poorest thermal noise performance or minimize its thickness. For example, with mirrors that use alternated layers of silica and titania-doped tantala, the goal is to reduce the total thickness of TiO2:Ta2O5 or replace it with higher-performing material (i.e., one with a higher refractive index, and/or lower loss angle) [4, 5]. Many other solutions are also being studied, including developing new thin films materials [6, 7], implementing crystalline coatings transferred to the silica substrate by optical contacting [8, 9], or even working at cryogenic temperature [10, 11]. To achieve the goal of reducing the coating thermal noise, the teams involved in this research have mostly continued to use ion beam sputtering as deposition process. However, alternative methods have also been considered, such as ion plating (IP), plasma-enhanced chemical vapor deposition (PECVD), or molecular beam epitaxy (MBE) for crystalline coatings. Like the others, the approach we present aims to reduce the thermal noise of layered materials while maintaining their reflectivity close to one. This is done by partially abandoning the use of alternating dielectric layers and replacing them with a metasurface.
Metasurfaces are planar devices structured at the wavelength scale. They have recently emerged as a new technology capable of pushing the boundaries of traditional optical components [12, 13], with manufacturing facilities now capable of structuring very large surfaces [16, 17]. Although dielectric resonant metasurfaces have theoretically demonstrated their potential as highly reflective coatings [14, 15], current demonstrations are still far from what is required for gravitational-waves detection. This is mainly because these resonant structures are highly sensitive to manufacturing errors, making every experimental realization difficult [18]. One way to mitigate this sensitivity is to combine the reflective metasurface with a multilayer stack. Dickmann and Kroker present such an assembly [19], in which a one-dimensional-structured metasurface and a multilayer Bragg mirror are placed on either side of a thick silica substrate [see Fig. 1.a]. This system functions as a Fabry-Perot interferometer, and its anti-resonant behavior is achieved by adjusting the optical thickness of the spacer through temperature control of the silica substrate.
Here, we propose an alternative approach [see Fig. 1.b]. First, a multilayer Bragg mirror is deposited on the substrate and covered by an additional low-index layer. Next, a high-index layer is deposited on the dielectric stack and 2D-structured to form the reflective metasurface. The additional low-index layer acts as a spacer between the Bragg mirror and the reflective metasurface, with its thickness optimized to ensure anti-resonant behavior.
We believe that our solution has advantages over the configuration proposed by Ref.[19] in terms of practical implementation within a Michelson interferometer dedicated to the detection of gravitational waves. First, the thermal tuning of the cavity must be compatible with the patterned heating of the mirror, which is already used to adjust the curvature radius of the interferometer’s end mirrors [20, 21] or stabilize the detector’s operation [22]. Second, the silica substrate would become part of the interferometer. This would make its manufacturing constraints much tighter in terms of rear face flatness and parallelism between the two faces. It would also be subject to the high laser fluence present in the interferometer’s arms. Third, using a 1D-structured metasurface makes the mirrors sensitive to the polarization state of the light. While not critical, this adds new constraints to the already complex design and operation of the instrument.
This paper aims to provide insight into the expected performance of the solution described by Fig.1.b, by examining different materials and design choices for the metasurface and the multilayer. The solutions studied here will be discussed and compared based on their expected reflectivity performance, manufacturing tolerances, and losses including both absorption and thermal noise.
II Metasurface design
First, we study the metasurface placed on top of the mirror independently. It consists of polarization-insensitive nano-cylinders with identical geometries. These elements are fabricated from a high-index dielectric material and are arranged on a silica substrate (), which is assumed to be semi-infinite during the metasurface design process. The design wavelength is set to nm, as it corresponds to several upcoming gravitational-wave detection projects [11].
As shown in Fig.2, for specific sets of geometrical parameters, these nano-cylinders can act as an array of dielectric resonators and behave as a reflective coating. When designing such structures, a higher refractive index facilitates the achievement of this resonant behavior. Thanks to its high refractive index ( at nm) and well-established structuring processes, amorphous silicon (a-Si) therefore appears as one of the most promising candidates for the metasurface material. In contrast, materials with a lower refractive index, such as tantala ( at nm), are less suitable for this type of resonant design, making it more difficult to achieve sufficiently high reflectivity (see Fig2c) . Considering a-Si as the pillars material, the periodicity of the metasurface is then optimized to ensure high reflectivity over the broadest possible range of parameters to increase fabrication tolerances. Fig.2 shows the effect of the periodicity on the metasurface reflectivity, evaluated using rigorous coupled wave analysis (RCWA), via Ansys Lumerical FDTD. After selecting a periodicity nm, the pillars diameter and height are chosen to lie at the center of the region exhibiting high reflectivity. The chosen set of parameters is represented by the white star in Fig.2b, respectively nm and nm.
Absorption in a-Si is not negligible and needs to be carefully evaluated, especially for applications such as gravitational wave detection, which requires incredibly low absorption levels (typically below the ppm). The extinction coefficient of amorphous silicon depends largely on the deposition process and varies widely in the recent literature, ranging from to at 1550 nm [25, 26, 27]. However, in the early 2000s, Domash et al. obtained much better results with hydrogenated amorphous silicon (a-Si:H) deposited by PECVD [28]. At the time, the goal was to produce narrow-band filters with thermally tunable central wavelengths. Experimental measurements confirmed extremely low absorption, although a-Si:H was used as the spacer material, with an extinction coefficient as low as at wavelength 1500 nm [29]. Additionally, recent studies on photovoltaic solar energy conversion have confirmed the feasibility of achieving such a low extinction coefficient for a-Si:H using PECVD as the deposition process [30]. Fig.2b shows the metasurface absorption for different extinction coefficient of a-Si, and highlights that the use of PECVD-deposited a-Si:H is a promising approach to limit absorption losses. In the case of an extinction coefficient of [29], the metasurface absorption is predicted around 4 ppm (see Fig.2b), while maintaining reflectance above .
III Association with a conventional multilayer Bragg mirror
III.1 Optimization of the additional low-index layer
Because of the resonant nature of the metasurface reflectivity, its performance remains sensitive to changes in the pillars geometry (see Fig.3b). It is therefore difficult to ensure the required performance over the entire range of typical manufacturing errors. The robustness of these structures can however be improved by coupling the metasurface with a multilayer stack [19], as described in Fig. 1.
In our approach shown in Fig. 1b, we achieve the anti-resonance behavior of the device by simply optimizing the thickness of the additional low-index layer which plays here the role of the spacer of the Fabry-Perot cavity formed by the two reflective structures (respectively the metasurface and the multilayer). The round-trip phase in the cavity is defined by [34]
where (respectively, ) is the phase change induced by the reflection on the metasurface (respectively, the Bragg mirror). To ensure anti-resonant behavior (i.e. maximum reflectivity), the round-trip phase must be an odd integer multiple of . This can be achieved with an infinite number of thicknesses, , that satisfy the following relation
All of these thicknesses are spaced by an interval of and depend on the phase terms, and , acquired at both sides of the cavity.
For simplicity, we will consider the multilayer to be an ideal Bragg mirror, with quarter-wave layers of alternating high and low index materials, such as TiO2:Ta2O5 ( at nm [32]) and silica. In such a quarter-wave multilayer stack, the phase change at reflection is equal to zero at the design wavelength [34], i.e. . This means that the cavity thicknesses that satisfy the condition (2) depend solely on the phase term acquired at the interface between the metasurface and the SiO2 layer. Consequently, the optimal thickness solutions are unaffected by the number of layers used for the Bragg mirror. The transmittance curves shown in Fig. 4 are therefore centered around the same peaks, whether the metasurface is combined with a 5-layer or an 11-layer Bragg mirror.
The solid lines represent the simulated transmittance obtained through RCWA simulations of the entire assembly. They exhibit a similar trend to the dotted lines, which are predicted by a classical Fabry-Perot model [34]:
However, deviations arise for small cavity lengths. We believe this is mainly because the metasurface reflection coefficient is obtained with the structured layer placed on a semi-infinite SiO2 substrate. In the assembly, however, it is placed on a silica layer, which makes the Fabry-Perot model of Eq. (3) inaccurate. As the cavity length increases, the semi-infinite substrate hypothesis used to optimize the metasurface becomes valid quickly, and the assembly response converges to the Fabry-Perot model.
Since smaller cavity thicknesses are preferable in terms of thermal noise, we select the first solution of (2), for which convergence to the Fabry-Perot model is ensured, namely nm. Because this solution is at the center of the stable plateaus predicted by the Fabry–Perot model, it allows for greater stability in the event of potential thickness errors when depositing the additional silica layer.
III.2 Impact on the metasurface tolerances
The manufacturing tolerances of our assembly need to be evaluated, since they are one of the major limitations of using a resonant reflective metasurface as a high-performance mirror solution. Figure 5 shows the performance of the reflective structure when slight changes are made to the geometrical parameters of the metasurface, and . This analysis is shown either for a single a-Si:H metasurface, and when coupled with an 11-layer TiO2:Ta2O5/SiO2 Bragg mirror. The white dashed lines in Fig.5 show the area where the transmittance is below 5 ppm. Although some parameters locally fulfill these conditions for the single metasurface case (Fig. 5a), this area is extremely small and is not represented for the sake of readability. However, as the number of layers used in the Bragg mirror increases, this dotted area expands, highlighting how coupling with the multilayer increases the metasurface manufacturing tolerances. In our case, only 11 layers are sufficient to relax the manufacturing tolerances, allowing for pillar diameter variations of up to nm while still meeting the required performance, making fabrication significantly more realistic.
Variations in transmittance still arise close to the resonance. Adding the Bragg mirror homogeneously lowers the transmittance values of the single metasurface, which facilitates performance that fulfills a certain threshold, such as ppm on a broad range of structure parameters. However, targeting a specific transmittance value, such as ppm, is more challenging as the metasurface response varies rapidly when close to the resonance. In that case, it is preferable to work slightly further from the resonance, to ensure a more stable behavior. However, this results in lower metasurface reflectivity and the need for more layers in the multilayer stack to ensure the required performance.
| Hybrid solution | Standard Bragg multilayer (37 layers) | |
| Total coating thickness | m | m |
| TiO2:Ta2O5 total thickness | m (6 quarter-wave layers) | m (19 quarter-wave layers) |
III.3 Thermal noise considerations
The final mirror design therefore combines an aSi:H metasurface with an 11-layer TiO2:Ta2O5/SiO2 Bragg mirror, with a nm-thick silica spacer between them. The final coating thickness is reported in Tab.1, together with that of a standard Bragg multilayer achieving the same level of reflectivity and requiring 39 quarter-wave alternated layers of TiO2:Ta2O5 and SiO2. Since thermal noise is closely related to the coating thickness, the reduction provided by such a hybrid mirror solution is very promising. In particular, the total thickness of TiO2:Ta2O5, which exhibits the poorest thermal-noise performance, is reduced by more than a factor of three.
A detailed thermal noise analysis of the entire design, and especially of the a-Si:H metasurface, has not yet been carried out. However, recent studies have highlighted the promising thermal noise performance of designs incorporating such structured layers [19, 33]. More generally, we believe that PECVD-deposited a-Si:H is an interesting material for developing high-performance mirrors with reduced thermal noise. One could even imagine using it in a multimaterial coating design [5, 25], with a-Si:H/Si3N4 pairs [28] replacing some of the bottom layers of the previously studied 11-layer TiO2:Ta2O5/SiO2 stack, thereby further reducing the overall coating thickness.
IV Conclusion
In conclusion, we investigated the feasibility of combining a resonant a-Si:H metasurface and a multilayer Bragg mirror. These two reflective structures are coupled through an additional silica layer, whose thickness is optimized to ensure maximum reflectivity. This hybrid approach significantly improves the robustness of the metasurface reflectivity to fabrication errors, while requiring only an 11-layer Bragg mirror to sufficiently relax fabrication tolerances, paving the way toward more accessible experimental validations. The proposed mirror solution can theoretically achieve the required performance for gravitational wave detection applications, while significantly reducing the coating thickness, making it a promising solution to potentially reduce the thermal noise as well.
All of the reflectivity and noise performance described here still need to be tested experimentally. This process may present unexpected challenges, as highlighted by the scattered light issues encountered by Ref. [36]. Although experimental validation is still required, we strongly believe that such hybrid designs are promising enough to encourage further investigations.
Acknowledgements.
Acknowledgments This work has been achieved thanks to the support of the French Centre National d’Etudes Spatiales - CNES and the company CILAS in the context of the joint laboratory LabTOP, the Virgo collaboration, A*MIDEX foundation for the funding of GRAVITERM project (AMX-22-RE-AB-091), and Agence Nationale de la Recherche for the funding of X-LOSM project (ANR-23-CE08-0027).References
- [1] B. P. Abbott and the LIGO Scientific Collaboration, ”Observation of Gravitational Waves from a Binary Black Hole Merger,” Phys. Rev. Lett. 116, 061102 (2016).
- [2] J. Degallaix, C. Michel, B. Sassolas, A. Allocca, G. Cagnoli, L. Balzarini, V. Dolique, R. Flaminio, D. Forest, M. Granata, B. Lagrange, N. Straniero, J. Teillon, and L. Pinard, ”TLarge and extremely low loss: the unique challenges of gravitational wave mirrors,” J. Opt. Soc. Am. A 36, C85-C94 (2019).
- [3] G.M. Harry, M.R. Abernathy, A.E. Becerra-Toledo, H. Armandula, E. Black, K. Dooley, M. Eichenfield, C. Nwabugwu, A. Villar, D.R.M. Crooks, G. Cagnoli, J. Hough, C.R. How, I. MacLaren, P. Murray, S. Reid, S. Rowan, P.H. Sneddon, M.M. Fejer, R. Route, S.D. Penn, P. Ganau, J. Mackowski, C. Michel, L. Pinard and A. Remillieu., “Titania-doped Tantala/Silica coatings for gravitational-wave detection,” Class. Quantum Grav. 24, 40 (2006).
- [4] M. Principe, ”Reflective coating optimization for interferometric detectors of gravitational waves,” Opt. Express 23, 10938-10956 (2015).
- [5] S. C. Tait, J. Steinlechner, M. M. Kinley-Hanlon, P. G. Murray, J. Hough, G. McGhee, F. Pein, S. Rowan, R. Schnabel, C. Smith, L. Terkowski, and I. W. Martin, ”Demonstration of the Multimaterial Coating Concept to Reduce Thermal Noise in Gravitational-Wave Detectors,” Phys. Rev. Lett. 125, 011102 (2020).
- [6] G. I. McGhee, V. Spagnuolo, N. Demos, S. C. Tait, P. G. Murray, M. Chicoine, P. Dabadie, S. Gras, J. Hough, G. A. Iandolo, R. Johnston, V. Martinez, O. Patane, S. Rowan, F. Schiettekatte, J. R. Smith, L. Terkowski, L. Zhang, M. Evans, I. W. Martin, and J. Steinlechner, ”Titania Mixed with Silica: A Low Thermal-Noise Coating Material for Gravitational-Wave Detectors,” Phys. Rev. Lett. 131, 171401 (2023).
- [7] A. Davenport, G. Vajente, N. Demos, A. Markosyan, S. Tait, M. M. Fejer, S. Gras, M. Evans, and C. S. Menoni, ”Minimization of Coating Thermal Noise in TiO2:GeO2 and SiO2 High Reflector Stacks for Gravitational Wave Detectors,” in Optica OIC — Optical Interference Coatings Conference 2025, Technical Digest Series (Optica Publishing Group, 2025), paper FA.4.
- [8] G. D. Cole, W. Zhang, M. Martin, J. Ye, M. Aspelmeyer, “Tenfold reduction of Brownian noise in high-reflectivity optical coatings,” Nat. Photonics 7, 644–650 (2013).
- [9] E. M. Gretarsson, A. M. Gretarsson, A. Juston, G. D. Cole, G. M. Harry, S. D. Penn, and B. N. Meagher, ”GaAs/Al0.92Ga0.08 As Crystalline Coatings in Gravitational Wave Detectors,” in Optica OIC — Optical Interference Coatings Conference 2025, Technical Digest Series (Optica Publishing Group, 2025), paper FA.6.
- [10] X. Koroveshi, L. Busch, E.Majorana, P.Puppo, P.Rapagnani, F.Ricci, P.Ruggi, S.Grohmann, ”Cryogenic payloads for the Einstein Telescope: Baseline design with heat extraction, suspension thermal noise modeling, and sensitivity analyses,” Phys. Rev. D 108, 123009 (2023).
- [11] T. Akutsu and the Scientific KAGRA Collaboration, ”KAGRA: 2.5 generation interferometric gravitational wave detector,” Nat. Astron. 3, 35-40 (2019).
- [12] P. Genevet, F. Capasso, F. Aieta, M. Khorasaninejad, and R. Devlin, ”Recent advances in planar optics: from plasmonic to dielectric metasurfaces,” Optica 4, 139-152 (2017).
- [13] A. I. Kuznetsov, M. L. Brongersma, J. Yao, M. K. Chen, U. Levy, D. P. Tsai, N. I. Zheludev, A. Faraon, A. Arbabi, N. Yu, D. Chanda, K. B. Crozier, A. V. Kildishev, H. Wang, J. K.W. Yang, J. G. Valentine, P. Genevet, J. A. Fan, O. D. Miller, A. Majumdar, J. E. Fröch, D. Brady, F. Heide, A. Veeraraghavan, N. Engheta, A. Alù, A. Polman, H. A. Atwater, P. Thureja, R. Paniagua-Dominguez, S. T. Ha, A. I. Barreda, J. A. Schuller, I. Staude, G. Grinblat, Y. Kivshar, S. Peana, S. F. Yelin, A. Senichev, V. M. Shalaev, S. Saha, A. Boltasseva, J. Rho, D. K. Oh, J. Kim, J. Park, Robert D., R. A. Pala, ”Roadmap for Optical Metasurfaces,” ACS Photonics 11, 816-865 (2024).
- [14] F. Brückner, D. Friedrich, T. Clausnitzer, M. Britzger, Ol. Burmeister, K. Danzmann, E.-B. Kley, A. Tünnermann, and R. Schnabel, ”Realization of a Monolithic High-Reflectivity Cavity Mirror from a Single Silicon Crystal,” Phys. Rev. Lett. 104, 163903 (2010).
- [15] M. Matiushechkina, A. B. Evlyukhin, V. A. Zenin, M. Heurs, B. N. Chichkov, ”High-efficiency silicon metasurface mirror on a sapphire substrate,” Opt. Mater. 138, 113618 (2023).
- [16] U. D. Zeitner, M. Banasch, and M. Trost, ”Potential of E-beam lithography for micro- and nano-optics fabrication on large areas,” J. Micro/Nanopattern. Mater. Metrol. 22, 041405 (2023).
- [17] J.-S. Park, S. W. D. Lim, A. Amirzhan, H. Kang, K. Karrfalt, D. Kim, J. Leger, A. Urbas, M. Ossiander, Z. Li, F. Capasso, ”All-Glass 100 mm Diameter Visible Metalens for Imaging the Cosmos,” ACS Nano. 18, 3187-3198 (2024).
- [18] M. Matiushechkina, A. B. Evlyukhin, R. Malureanu, V. A. Zenin, T. Yezekyan, A. Lavrinenko, S. I. Bozhevolnyi, B. N. Chichkov, M. Heurs, ”Design and Experimental Demonstration of Wavelength-Selective Metamirrors on Sapphire Substrates,” Adv. Photon. Res. 6, 2400116 (2024).
- [19] J. Dickmann, S. Kroker, ”Highly reflective low-noise etalon-based meta-mirror,” Phys. Rev. D 98, 082003 (2018).
- [20] R. Lawrence, D. Ottaway, M. Zucker, and P. Fritschel, “Active correction of thermal lensing through external radiative thermal actuation,” Opt. Lett. 29, 2635–2637 (2004).
- [21] J. Degallaix, ”Compensation of strong thermal lensing in advanced interferometric gravitational waves detectors,” Ph.D. dissertation (University of Western Australia, 2006).
- [22] T. Hardwick, V. J. Hamedan, C. Blair, A. C. Green, and D. Vander-Hyde, ”Demonstration of dynamic thermal compensation for parametric instability suppression in Advanced LIGO,” Class. Quantum Grav. 37, 205021 (2020).
- [23] G. M. Harry and the LIGO Scientific Collaboration, ”Advanced LIGO: the next generation of gravitational wave detectors,” Class. Quantum Grav. 27, 084006 (2010).
- [24] M. Le Jean, J. Degallaix, D. Hofman, L. Pinard, D. Forest, M. Granata, C. Michel, J. Steinlechner, C. Amra, M. Lequime, and M. Zerrad, ”Comparison of arm cavity optical losses for the two wavelengths of the Einstein telescope gravitational wave detector,” Class. Quantum Grav. 41, 225004 (2024).
- [25] J. Steinlechner, I. W Martin, R. Bassiri, A. Bell, M. M. Fejer, J. Hough, A. Markosyan, R. K. Route, and S. Rowan, ”Optical Absorption of Ion-Beam Sputtered aSi Coatings,” Phys. Rev. D 93, 062005 (2016).
- [26] I. A. Birney, J. Steinlechner, Z. Tornasi, S. MacFoy, D. Vine, A. S. Bell, D. Gibson, J. Hough, S. Rowan, P. Sortais, S. Sproules, S. Tait, I. W. Martin, and S. Reid, ”Amorphous Silicon with Extremely Low Absorption: Beating Thermal Noise in Gravitational Astronomy,” Phys. Rev. Lett. 121, 191101 (2018).
- [27] M. Molina-Ruiz, A. Markosyan, R. Bassiri, M. M. Fejer, M. Abernathy, T. H. Metcalf, X. Liu, G. Vajente, A. Ananyeva, F. Hellman, ”Hydrogen-Induced Ultralow Optical Absorption and Mechanical Loss in Amorphous Silicon for Gravitational-Wave Detectors,” Phys. Rev. Lett. 131, 256902 (2023).
- [28] L. Domash, M. Wu, N. Nemchuk, and E. Ma, ”Tunable and Switchable Multiple-Cavity Thin Film Filters,” J. Lightwave Technol. 22, 126-135 (2004).
- [29] L. H. Domash, E. Ma, M. Lourie, W. Sharfin, and M. Wagner, ”Broadly tunable thin film interference coatings: active thin films for telecom applications,” in Optical Devices for Fiber Communication IV, M. J. F. Digonnet, ed., Proc. SPIE 4989, 161-167 (2003).
- [30] P. K. Tawalare, ”Optimizing photovoltaic conversion of solar energy,” AIP Advances 11, 100701 (2021).
- [31] C. Zhang, L. Chen, Z. Lin, J. Song, D. Wang, M. Li, O. Koksal, Z. Wang, G. Spektor, D. Carlson, H. J. Lezec, W. Zhu, S. Papp and A. Agrawal, ”Tantalum pentoxide: a new material platform for high-performance dielectric metasurface optics in the ultraviolet and visible region,” Light Sci. Appl. 13, 23 (2024).
- [32] M. Granata, A. Amato, L. Balzarini, M. Canepa, J. Degallaix, D. Forest, V. Dolique, L. Mereni, C. Michel, L. Pinard, B. Sassolas, J. Teillon, G. Cagnoli, ”Amorphous optical coatings of present gravitational-wave interferometers,” Class. Quantum Grav. 37, 095004 (2020).
- [33] J. Dickmann, C. B. Rojas Hurtado, R. Nawrodt, S. Kroker, ”Influence of polarization and material on Brownian thermal noise of binary grating reflectors,” Phys. Lett. A 382, 2275-2281 (2018).
- [34] C. Amra, M. Lequime, and M. Zerrad, Electromagnetic Optics of Thin-Film Coatings - Light Scattering, Giant Field Enhancements, and Planar Microcavities, Cambridge University Press (2021).
- [35] A. Patoux, G. Agez, C. Girard, V. Paillard, P. R. Wiecha, A. Lecestre, F. Carcenac, G. Larrieu, and A. Arbouet, ”Challenges in nanofabrication for efficient optical metasurfaces,” Sci. Rep. 11, 5620 (2021).
- [36] J. Dickmann, S. Sauer, J. Meyer, M. Gaedtke, T. Siefke, U. Brückner, J. Plentz, and S. Kroker, ”Experimental realization of a 12,000-finesse laser cavity based on a low-noise microstructured mirror,” Commun Phys 6, 16 (2023).